OpenStudy (anonymous):

A canoe can go 24 km downstream in 3 hours. The return trip takes 4 hours. What is the speed of the current? You will earn 2 points for writing the correct system of equations for this word problem. Remember, two equations must be given for the system of equations. You will earn 3 points for solving it correctly and answering the question in a complete sentence.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

if v1 is canoe velocity and v2 is the river velocity. When canoe goes down stream means in the direction of river, the net velocity is the sum of v1+v2. similarly in the upstream v1-v2. Distance is same in both cases i.e. 24 km. and time is in the down stream 3hrs and in the up stream 4 hrs. distance = velocity * time. down stream=> (v1+v2)* 3 = 24 up stream => (v1-v2)*4 = 24. By these equations we can easily solve for V1 and V2

OpenStudy (anonymous):

v2 is current speed

OpenStudy (ny,ny):

NY,NY 0 Let s= still water. Let c = current speed. t = time. D = distance. To have the canoe travel with the current (downstream), (s+c) t = D To have the canoe travel against the current (upstream), (s-c) t = D ---- (s+c) t = D (s+c) (3) = 24 (s+c) = 8 (s-c) t = D (s-c) (4) = 24 ----- (s-c) = 6 + (s+c) = 8 ------------ 2s + 0 = 14 s=7 ---- (s+c) = 8 ((7)+c) = 8 c=1 did you understand that ^^ ?

OpenStudy (anonymous):

yes thank you

OpenStudy (ny,ny):

:)

OpenStudy (anonymous):

No